OpenStudy (anonymous):

Hi :) Please help refresh my memory :0 Find the slope and the intercepts of: g(x)=54x-1

OpenStudy (anonymous):

given y=54x-1 it is of the form y=mx +c hence slope is 54 and the y-intercept is -1 ( f(x) crosses y-axis at (0,-1)) for find ing where f(x) crosses x-axis we put y or(f(x)) =0 hence here 0=54x-1 or x=1/54 hence f(x) crosses x-axis at (1/54,0)

OpenStudy (anonymous):

Thanks so much!

OpenStudy (anonymous):

welcome